OpenStudy (anonymous):

In a group of 200 high school students, 36 are taking biology, 52 are taking Spanish, and 126 are taking neither biology nor Spanish. If one of these 200 students is to be chosen at random, what is the probability that the student chosen is taking biology but not Spanish? A. 7% B. 11% C. 18% D. 37%

OpenStudy (anonymous):

I think its B idk

OpenStudy (anonymous):

I'd say B.

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

11%

OpenStudy (nathan917):

126 (neither Bio. nor Spanish) +52 (Spanish) ___ 178 (total of students that can't be chosen) 200 (total students) -178 (total it can't be chosen) ____ 22 22/200 = 11/100 = 11%

OpenStudy (nathan917):

You were correct young. Metal for you :)